# Optimizing reaction and transport fluxes in temperature gradient-driven chemical reaction-diffusion systems

### v1

@@ II.3.1 Model description
The spatial extension of the system was considered by introducing the chemical diffusion of all compounds. We limited our study to closed one-dimensional systems. A stationary temperature gradient $\theta(\bar{x})$ was imposed over the entire system. This corresponds to an unstirred system, with free internal diffusion of chemical compounds, without matter exchange with the surroundings. The system was described using a partial differential equation set (see Eqs. (25aibh)), which was numerically solved for a large array of parameters, focusing on determining its steady-state (see Appendix A). The characteristic parameter $l_{0}$ must be introduced to account for the spatial extension, describing a one dimensional system:
@@
Chemical diffusion is described based on the diffusion parameter $d_{i}$, which is nondimensionalized by the characteristic parameter $D_{0}$ corresponding to the average diffusion constant of the chemical reactants. Consistently with the description of the chemical exchange kinetics of the previous model, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$, and the exchange ratio $\lambda$ (see Eqs. (25ag)). The characteristic time $t_{0}$ is additionally fixed by $l_{0}$ and $D_{0}$ (see Table. 1). All parameters corresponding to the chemical reactions are defined as in the stirred-compartment model; however, they are now functions of position $\bar{x}$.
@@
  • We neglected the influence of temperature on the diffusion coefficient. A few simulations were performed, indicating that considering the variability of $d_{i}$ with $\theta$ did not qualitatively change the system behavior; this essentially led to an increase in the asymmetry between the cold and warm regions (see Fig. S2 in SI).
  • The Soret effect was neglected, which is expected to be satisfactory for systems based on small molecular reactants.[37]
  • Convection was not considered in this study. This assumption is anticipated to be satisfactory for sufficiently small systems in the direction of gravity, or sufficiently highly viscous systems, so that the major transport mode is processed by molecular diffusion. This condition is compatible with a large system in which the temperature gradient is applied on the horizontal $x$ axis, as long as the vertical spatial extension is sufficiently small (e.g., for horizontal capillary systems).
@@
II.3.2 Steady-state flux profiles
@@
In the spatial region that is sufficiently close to the $\bar{x}=0.5$ cold/warm boundary—i.e. at a scale smaller than the reaction-diffusion length $w$—the chemical diffusion is much faster than chemical reactions. The concentrations are consequently essentially constant on each side of the temperature boundary; the system is then equivalent to the “two compartments” model in the vicinity of the boundary. The values of the flux $\varphi$ and force $\alpha$ determined for the two-compartments model correspond to the values observed at the boundary $x=0.5^{+}$ and $x=0.5^{-}$.
When $w>0.5$, the reaction-diffusion length is larger than the sizes of the cold and warm regions $\bar{x}\in[0,0.5]$ and $\bar{x}\in[0.5,1]$. Both are homogeneous; the full system is then equivalent to the stirred compartment model. This system is characterized by fast diffusion throughout the system, while the chemical reaction flux and dissipation are evenly distributed throughout the entire system (see Fig. 3B).
When $w<0.5$, an exponential decrease in $\varphi$ and $\alpha$ values is observed, from their boundary values at $\bar{x}=0.5$ to either $x=0$ or $x=1$, with an half-length equal to $w$. If $w\ll 1$, the system activity is restricted to a short region close to the boundary, most of the system being at equilibrium and characterized by $\varphi\approx 0$, $\alpha\approx 0$, $\sigma\approx 0$ (see Fig. 3D).
@@
The spatial extension of chemical activity in each region becomes asymmetric when the temperature gradient intensity is sufficiently high. This can be described by considering the difference in the kinetic rate in each region based on Eq. (25aea)-(25aeb), as:
@@
When $\beta\theta_{g}>1$, the system becomes highly nonlinear, with a large behavioral asymmetry between each region. Typically, situations with $w_{\text{cold}}>1$ and $w_{\text{warm}}\ll 1$ can be obtained, with a system homogeneous in the cold zone, and restricted to a short region in the warm zone (see Fig. 3E).
@@
  • Fast-diffusion systems (i.e. $w\gg 1$, see Fig. 3G) are characterized by homogeneous concentration profiles. This induces a linear variation in $\alpha$ and $\varphi$, reproducing the temperature profile shape, and thus a parabolic shape for $\sigma=\alpha\varphi$. Consequently, the system is more active at the system extremities, with $\alpha=0$, $\varphi=0$ and $\sigma=0$ close to $\bar{x}=\nicefrac{{1}}{{2}}$.
  • In fast-reaction systems (i.e. $w\ll 1$, see Fig. 3I), the central inactive region characterized by $\alpha\approx 0$, $\varphi\approx 0$ and $\sigma\approx 0$ becomes much larger. The chemical reaction flux is restricted at each system extremity, on a region of length $w_{\text{cold}}$ at $\bar{x}\approx 0$ and $w_{\text{warm}}$ at $\bar{x}\approx 1$. In turn, chemical transport spreads, linking these two extremities.
This general behavior implies a de facto spatial separation of the different chemical processes. Three regions can be identified: cold and warm regions at each extremity, of half-lengths equal to $w_{\text{cold}}$ and $w_{\text{warm}}$, separated by a central region of length approximately equal to $1-2w$. The cold and warm regions are dominated by chemical reactions ($\sigma\approx\sigma_{r}$), and the central region is dominated by chemical transport ($\sigma\approx\sigma_{d}$). A spatially extended reaction-diffusion cycle is thus set in motion; the compounds are transported across the entire system between each chemically active region—the central transport region shrinking with $w$ (see Fig. 3F). This is in sharp contrast with the Heaviside gradient, where both chemical reactions and transport are effective at the same spatial locations, close to the temperature boundary (see Fig. 3A).
A direct consequence of this chemical transport between spatially separated regions is a sharp loss of efficiency of the dissipative processes for small $w$: the boundary values of $\alpha$ and $\varphi$ are not anymore scale-independent as observed with Heaviside profiles, but decrease proportionally to $w$ when $w\ll 1$. Thus, slow diffusion systems (or large systems) are much closer to equilibrium for linear temperature gradients than for Heaviside gradients.
Finally, the high profile asymmetrization induced by high $\beta\theta_{g}$ values leads to non-trivial and complex flux profiles (see Fig. 3J). The central region around $\bar{x}=\nicefrac{{1}}{{2}}$ is characterized by a non-negligible positive flux, corresponding to a low-activity extension of the warm region. Consequently, the spatial position of the zero flux is localized at $\bar{x}<\nicefrac{{1}}{{2}}$; the cold region of the negative flux shrinks, and the warm region of the positive flux stretches.
@@
A last temperature gradient, was defined, as a mix between a linear and a Heaviside gradient. This is defined as $\theta=-\theta_{g}$ for $\bar{x}<\bar{x}_{1}=\nicefrac{{1}}{{4}}$, $\theta=+\theta_{g}$ for $\bar{x}>\bar{x}_{2}=\nicefrac{{3}}{{4}}$, and a linear variation of $\theta$ from $-\theta_{g}$ to $+\theta_{g}$ for $\bar{x}\in[\nicefrac{{1}}{{4}},\nicefrac{{3}}{{4}}]$. This comes down to a system in contact with two spatially extended external thermostats.
@@
The behaviors of both the linear and Heaviside profiles can be recognized (see Fig. 4). Reaction-diffusion cycles are now set in motion, with two chemical reaction zones localized at $\bar{x}_{1}$ and $\bar{x}_{2}$, connected by an active chemical transport region across $\bar{x}\in[\bar{x}_{1},\bar{x}_{2}]$.
It appears that controlling the precise shape of an imposed temperature gradient enables precise control of the shape of the induced reaction-diffusion cycle. Steady dissipative chemical reactions can be localized in regions characterized by temperature inflection points (i.e. high values of $\nicefrac{{\partial^{2}T}}{{\partial x^{2}}}$), whereas chemical transport can be sustained between these active regions along temperature gradients (i.e. following high $\nicefrac{{\partial T}}{{\partial x}}$ values).
@@ II.3.3 General description
The steady-state can be described by a set of parameters $\alpha$, $\varphi$, $\sigma_{r}$, $\sigma_{d}$ and $\sigma_{t}=\sigma_{r}+\sigma_{d}$. Their maximal value assesses the optimal activity in a specific location, and their integral value assesses the optimal activity of the system as a whole (see Appendix D.2).
@@
In the limit cases $w\ll 1$ and $w\gg 1$, and when the system is close to equilibrium (i.e. $\beta\theta_{g}<1$), the ratio of each variable to its reference value is only dependent on the reaction-diffusion length $w$ as defined in Eq. (15). The corresponding values are summarized in Table 2 (see Fig. S3 in SI for graphical representation).
@@
The situation $w\gg 1$ (i.e., small systems) is essentially independent of the scale. This implies that, as long as the system dimension is smaller than the reaction-diffusion length, the steady-state behavior is essentially described by the characteristic values of Eqs. (18), modulated by a constant factor that depends on the shape of the temperature gradient profile. Diffusion dissipation is the only varying parameter; it is proportional to $w^{-2}$, and thus is essentially negligible for small systems. This situation is the optimal case for extracting energy from the temperature gradient (via $\sigma_{t}$ maximization), as well as driving nonequilibrium chemical reaction (via both $\sigma_{r}$ and $y_{r}$ maximization).
@@
When $w\ll 1$ (i.e. large systems), the system is only effective on a small fraction of the full spatial extension, except for the transport. This implies that all steady-state parameters are proportional to $w^{n}$, with $n\in[1,3]$, except for $\sigma_{d}$ which remains constant. This is the optimal situation for transporting chemical compounds over large distances.
@@
The intermediate case $w\sim 1$ is characterized by the maximal value of $\sigma_{r}$, i.e. by optimal nonequilibrium transport. In this regime, the chemical reaction is also important, with a 50% dissipation of entropy in both processes in the case $w=1$. This situation is thus optimal for observing nonequilibrium cycles of reaction-diffusion, as both diffusion and reaction can be simultaneously efficient.
@@
This behavior is further influenced by the system symmetry as defined by $\lambda$ and $\kappa$ parameters. Optimal nonequilibrium reaction fluxes are obtained for symmetrical systems characterized by $\kappa=1$ when $w>1$ (Fig. 5A), and $\kappa=1$ and $\lambda=1$ when $w<1$ (Fig. 5C and E). As observed by Liang et al,[34] an asymmetry in the diffusion coefficient, characterized by $\lambda\neq 1$, can lead to so-called “emergent thermophoretic effect”, implying that the coupling of the chemical reaction to the temperature gradient can lead to concentrating the reactants in either the colder or the warmer region.
@@
This effect is inefficient in small systems (Fig. 5B). Thus, a sufficiently large system is required (Fig. 5D and F) such that the dissipative processes are dominated by chemical transports. This process is especially efficient for values of $\lambda\neq 1$, and needs to be driven away from the equilibrium by high intensity gradients (Fig. 5F). Optimal $\kappa$ values are additionally observed, with $\kappa>1$ values associated with $\lambda<1$ (and consequently $\kappa<1$ values associated with $\lambda>1$ values).
@@ II.3.4 System scaling
From a general perspective, the efficiency of the system to be driven far from the equilibrium state is directly linked to the maximization of $\beta$, $\gamma$, $\theta_{g}$ and $\rho$. This trivially implies that the best chemical reaction candidates are characterized by high activation energies, high reaction enthalpies, high reaction rates, and should be placed in high temperature gradients.
More importantly, the results specifically emphasize the necessity of adapting the spatial scale of the chemical setup to the characteristic scale of the reaction-diffusion system. The temperature gradient should also be chosen carefully, so that the system is as symmetric as possible at the mean temperature, in terms of chemical diffusion (i.e. $D_{1}(T_{0})\approx D_{2}(T_{0})$) and balance between the reactants (i.e. $K(T_{0})\approx 1$).
@@
The system spatial dimension $l_{0}$ should be adjusted accordingly. Optimal energy extraction and nonequilibrium chemical reactions are obtained for $l_{0}<\mathcal{L}_{0}$, optimal chemical transport over large distances for $l_{0}>\mathcal{L}_{0}$, and optimal reaction-diffusion cycle for $l_{0}\sim\mathcal{L}_{0}$,
@@
A reaction of reversible peptide exchange mediated by N-methyl-cysteine[38] is a good candidate for being driven by temperature gradients. It possesses high activation energies, and its equilibrium constant is $K=1$ close to the ambient temperature:[39]
@@
This is a second order reaction. An apparent first order constant (that depends on the global reactant concentrations) can be evaluated from $k_{0}^{+}$ to $k_{0,\text{app}}=1.4\times 10^{-6}$ s-1 assuming a concentration $c_{0}=10^{-3}$ M.
The diffusion coefficients of short oligopeptides[40, 41] are approximately equal to $D_{0}=5\times 10^{-10}$ m2.s-1, and will be assumed to be identical (i.e. $\lambda=1$).
In that conditions, the critical parameters of this reaction-diffusion system are $\mathcal{L}_{0}=1.3$ cm and $\mathcal{T}_{0}=17$C. In order to optimize this chemical system for efficient reaction-diffusion cycles, an optimal setup should be on the centimeter scale, with a mean temperature of $40$C. Nonlinear behavior should be obtained when subjected to a gradient of at least ±$\pm 17$C.
@@
  • $l_{1}=3$ cm: $w_{\text{cold}}=1.27$, $w_{\text{warm}}=0.19$. This corresponds to an intermediate system in which the activity is fully extended in the cold region, and restricted in the region close to the warm boundary, dominated by a dissipative reaction-diffusion cycle across both regions.

### v2

@@ II.3.1 Model description
The spatial extension of the system was considered by introducing the chemical diffusion of all compounds. We limited our study to closed one-dimensional systems. A stationary temperature gradient $\theta(\bar{x})$ was imposed over the entire system. This corresponds to an unstirred system, with free internal diffusion of chemical compounds, without matter exchange with the surroundings. The system was described using a set of partial differential equations (see Eqs. (25aibh)), which was numerically solved for a wide range of parameters, with a focus on determining its steady state (see Appendix A). The characteristic parameter $l_{0}$ must be introduced to account for the spatial extension, describing a one dimensional system:
@@
Chemical diffusion is described based on the diffusion parameter $d_{i}$, which is nondimensionalized by the characteristic parameter $D_{0}$ corresponding to the average diffusion constant of the chemical reactants. Consistently with the description of the chemical exchange kinetics of the previous model, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$ and the exchange ratio $\lambda$ (see Eqs. (25ag)). The characteristic time $t_{0}$ is additionally fixed by $l_{0}$ and $D_{0}$ (see Table. 1). All parameters corresponding to chemical reactions are defined as in the stirred compartment model; however, they are now functions of position $\bar{x}$.
@@
  • We neglected the influence of temperature on the diffusion coefficient. Several simulations were performed, indicating that considering the variability of $d_{i}$ with $\theta$ did not qualitatively change the system behavior; this essentially led to an increase in the asymmetry between the cold and warm regions (see Fig. S2 in SI).
  • The Soret effect was neglected, which is expected to be satisfactory for systems based on small molecular reactants[37].
  • Convection was not considered in this study. This assumption is anticipated to be satisfactory for sufficiently small systems in the direction of gravity or sufficiently highly viscous systems, so that the major transport mode is processed by molecular diffusion. This condition is compatible with a large system in which the temperature gradient is applied on the horizontal $x$ axis, as long as the vertical spatial extension is sufficiently small (e.g., for horizontal capillary systems).
@@
II.3.2 steady state flux profiles
@@
In the spatial region that is sufficiently close to the $\bar{x}=0.5$ cold / warm boundary—that is, at a scale smaller than the reaction-diffusion length $w$—chemical diffusion is much faster than chemical reactions. The concentrations are consequently essentially constant on each side of the temperature boundary; the system is then equivalent to the “two compartments” model in the vicinity of the boundary. The values of the flux $\varphi$ and force $\alpha$ determined for the two-compartment model correspond to the values observed at the boundaries $x=0.5^{+}$ and $x=0.5^{-}$.
When $w>0.5$, the reaction-diffusion length is larger than the sizes of the cold and warm regions $\bar{x}\in[0,0.5]$ and $\bar{x}\in[0.5,1]$. Both are homogeneous; the full system is then equivalent to the stirred-compartment model. This system is characterized by fast diffusion throughout the system, whereas the chemical reaction flux and dissipation are evenly distributed throughout the entire system (see Fig. 3B).
When $w<0.5$, an exponential decrease is observed in the values of $\varphi$ and $\alpha$, from their boundary values at $\bar{x}=0.5$ to $x=0$ or $x=1$, with a half-length equal to $w$. If $w\ll 1$, the activity of the system is restricted to a short region close to the boundary, most of the system being in equilibrium and characterized by $\varphi\approx 0$, $\alpha\approx 0$, $\sigma\approx 0$ (see Fig. 3D).
@@
The spatial extension of chemical activity in each region becomes asymmetric when the intensity of the temperature gradient is sufficiently high. This can be described by considering the difference in the kinetic rate in each region based on Eq. (25aea)-(25aeb), as:
@@
When $\beta\theta_{g}>1$, the system becomes highly nonlinear, with a large behavioral asymmetry between each region. Typically, situations with $w_{\text{cold}}>1$ and $w_{\text{warm}}\ll 1$ can be obtained, with a homogeneous system in the cold zone and restricted to a short region in the warm zone (see Fig. 3E).
@@
  • Fast-diffusion systems (i.e. $w\gg 1$, see Fig. 3G) are characterized by homogeneous concentration profiles. This induces a linear variation in $\alpha$ and $\varphi$, reproducing the shape of the temperature profile, and thus a parabolic shape for $\sigma=\alpha\varphi$. Consequently, the system is more active at the extremities of the system, with $\alpha=0$, $\varphi=0$ and $\sigma=0$ close to $\bar{x}=\nicefrac{{1}}{{2}}$.
  • In fast reaction systems (i.e. $w\ll 1$, see Fig. 3I), the central inactive region characterized by $\alpha\approx 0$, $\varphi\approx 0$ and $\sigma\approx 0$ becomes much larger. The chemical reaction flux is restricted at each extremity of the system, in a region of length $w_{\text{cold}}$ at $\bar{x}\approx 0$ and $w_{\text{warm}}$ at $\bar{x}\approx 1$. In turn, chemical transport spreads, linking these two extremities.
This general behavior implies a spatial separation de facto of the different chemical processes. Three regions can be identified: cold and warm regions at each extremity, of half-lengths equal to $w_{\text{cold}}$ and $w_{\text{warm}}$, separated by a central region of length approximately equal to $1-2w$. The cold and warm regions are dominated by chemical reactions ($\sigma\approx\sigma_{r}$), and the central region is dominated by chemical transport ($\sigma\approx\sigma_{d}$). Thus, a spatially extended reaction-diffusion cycle is set in motion; the compounds are transported across the entire system between each chemically active region, the central transport region shrinking with $w$ (see Fig. 3F). This is in sharp contrast with the Heaviside gradient, where both chemical reactions and transport are effective at the same spatial locations, close to the temperature boundary (see Fig. 3A).
A direct consequence of this chemical transport between spatially separated regions is a sharp loss of efficiency of the dissipative processes for small $w$: the boundary values of $\alpha$ and $\varphi$ are no longer scale independent as observed with Heaviside profiles, but decrease proportionally to $w$ when $w\ll 1$. Thus, slow-diffusion systems (or large systems) are much closer to equilibrium for linear temperature gradients than for Heaviside gradients.
Finally, the high asymmetrization of the profile induced by high $\beta\theta_{g}$ values leads to nontrivial and complex flux profiles (see Fig. 3J). The central region around $\bar{x}=\nicefrac{{1}}{{2}}$ is characterized by a non-negligible positive flux, corresponding to a low-activity extension of the warm region. Consequently, the spatial position of the zero flux is localized at $\bar{x}<\nicefrac{{1}}{{2}}$; the cold region of the negative flux shrinks and the warm region of the positive flux stretches.
@@
A last temperature gradient was defined as a mix between a linear and a Heaviside gradient. This is defined as $\theta=-\theta_{g}$ for $\bar{x}<\bar{x}_{1}=\nicefrac{{1}}{{4}}$, $\theta=+\theta_{g}$ for $\bar{x}>\bar{x}_{2}=\nicefrac{{3}}{{4}}$, and a linear variation of $\theta$ from $-\theta_{g}$ to $+\theta_{g}$ for $\bar{x}\in[\nicefrac{{1}}{{4}},\nicefrac{{3}}{{4}}]$. This comes down to a system in contact with two spatially extended external thermostats.
@@
The behaviors of both the linear and Heaviside profiles can be recognized (see Fig. 4). The reaction-diffusion cycles are now in motion, with two chemical reaction zones located at $\bar{x}_{1}$ and $\bar{x}_{2}$, connected by an active chemical transport region across $\bar{x}\in[\bar{x}_{1},\bar{x}_{2}]$.
It appears that controlling the precise shape of an imposed temperature gradient enables precise control of the shape of the induced reaction-diffusion cycle. Steady dissipative chemical reactions can be localized in regions characterized by temperature inflection points (i.e., high values of $\nicefrac{{\partial^{2}T}}{{\partial x^{2}}}$), whereas chemical transport can be sustained between these active regions along temperature gradients (i.e., following high values of $\nicefrac{{\partial T}}{{\partial x}}$.
@@ II.3.3 General description
The steady state can be described by a set of parameters $\alpha$, $\varphi$, $\sigma_{r}$, $\sigma_{d}$ and $\sigma_{t}=\sigma_{r}+\sigma_{d}$. Their maximal value assesses the optimal activity in a specific location, and their integral value assesses the optimal activity of the system as a whole (see the appendix D.2).
@@
In the limit cases $w\ll 1$ and $w\gg 1$, and when the system is close to equilibrium (i.e. $\beta\theta_{g}<1$), the ratio of each variable to its reference value is only dependent on the reaction-diffusion length $w$ as defined in Eq. (15). The corresponding values are summarized in Table 2 (see Fig. S3 in SI for a graphical representation).
@@
The situation $w\gg 1$ (that is, small systems) is essentially independent of scale. This implies that, as long as the dimension of the system is smaller than the length of the reaction-diffusion, the steady-state behavior is essentially described by the characteristic values of Eqs. (18), modulated by a constant factor that depends on the shape of the temperature gradient profile. Diffusion dissipation is the only varying parameter; it is proportional to $w^{-2}$, and thus is essentially negligible for small systems. This situation is the optimal case for extracting energy from the temperature gradient (by maximizing $\sigma_{t}$, as well as driving nonequilibrium chemical reaction (c=by maximizing both $\sigma_{r}$ and $y_{r}$).
@@
When $w\ll 1$ (that is, large systems), the system is effective only on a small fraction of the full spatial extension, except for transport. This implies that all steady-state parameters are proportional to $w^{n}$, with $n\in[1,3]$, except for $\sigma_{d}$, which remains constant. This is an optimal situation for transporting chemical compounds over large distances.
@@
The intermediate case $w\sim 1$ is characterized by the maximal value of $\sigma_{r}$, that is, by optimal nonequilibrium transport. In this regime, the chemical reaction is also important, with a dissipation of entropy equally shared in both processes for $w=1$. This situation is thus optimal for observing nonequilibrium cycles of reaction-diffusion, as both diffusion and reaction can be simultaneously efficient.
@@
This behavior is also influenced by the symmetry of the system defined by the parameters $\lambda$ and $\kappa$. Optimal nonequilibrium reaction fluxes are obtained for symmetrical systems characterized by $\kappa=1$ when $w>1$ (Fig. 5A), and $\kappa=1$ and $\lambda=1$ when $w<1$ (Fig. 5C and E). As observed by Liang et al[34], an asymmetry in the diffusion coefficient, characterized by $\lambda\neq 1$, can lead to so-called “emergent thermophoretic effect”, implying that the coupling of the chemical reaction to the temperature gradient can lead to concentrating the reactants in either the colder or the warmer region.
@@
This effect is inefficient in small systems (Fig. 5B). Thus, a sufficiently large system is required (Fig. 5D and F) such that the dissipative processes are dominated by chemical transports. This process is especially efficient for values of $\lambda\neq 1$, and needs to be driven away from the equilibrium by high intensity gradients (Fig. 5F). Optimal $\kappa$ values are also observed, with $\kappa>1$ values associated with $\lambda<1$ (and consequently $\kappa<1$ values associated with $\lambda>1$ values).
@@ II.3.4 System scaling
From a general perspective, the efficiency of the system to be driven far from the equilibrium state is directly linked to the maximization of $\beta$, $\gamma$, $\theta_{g}$ and $\rho$. This trivially implies that the best chemical reaction candidates are characterized by high activation energies, high reaction enthalpies, and high reaction rates and should be placed in high-temperature gradients.
More importantly, the results specifically emphasize the necessity of adapting the spatial scale of the chemical setup to the characteristic scale of the reaction-diffusion system. The temperature gradient should also be chosen carefully, so that the system is as symmetric as possible at the mean temperature, in terms of chemical diffusion (that is, $D_{1}(T_{0})\approx D_{2}(T_{0})$) and balance between the reactants (i.e. $K(T_{0})\approx 1$).
@@
The system spatial dimension $l_{0}$ should be adjusted accordingly. Optimal energy extraction and nonequilibrium chemical reactions are obtained for $l_{0}<\mathcal{L}_{0}$, optimal chemical transport over large distances for $l_{0}>\mathcal{L}_{0}$, and the optimal reaction-diffusion cycle for $l_{0}\sim\mathcal{L}_{0}$,
@@
Reversible peptide exchange reaction mediated by N-methylcysteine is a good candidate for being driven by temperature gradients[38]. It possesses high activation energies, and its equilibrium constant is $K=1$ close to the ambient temperature[39]:
@@
This is a second-order reaction. An apparent first-order constant (that depends on the global reactant concentrations) can be evaluated from $k_{0}^{+}$ to $k_{0,\text{app}}=1.4\times 10^{-6}$ s-1 assuming a concentration $c_{0}=10^{-3}$ M.
The diffusion coefficients of short oligopeptides[40, 41] are approximately equal to $D_{0}=5\times 10^{-10}$ m2.s-1, and will be assumed that they are identical (that is, $\lambda=1$).
Under such conditions, the critical parameters of this reaction-diffusion system are $\mathcal{L}_{0}=1.3$ cm and $\mathcal{T}_{0}=17$C. In order to optimize this chemical system for efficient reaction-diffusion cycles, an optimal setup should be on the centimeter scale, with a mean temperature of $40$C. Nonlinear behavior should be obtained when subjected to a gradient of at least ±$\pm 17$C.
@@
  • $l_{1}=3$ cm: $w_{\text{cold}}=1.27$, $w_{\text{warm}}=0.19$. This corresponds to an intermediate system in which the activity is fully extended in the cold region and restricted in the region close to the warm boundary, dominated by a dissipative reaction-diffusion cycle across both regions.